Circular No. 6277

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

HD 49798 AND 2E 0050.1–7247 G. L. Israel, International School for Advanced Studies, Trieste; L.

Stella, Astronomical Observatory of Rome; L. Angelini and N. E. White, Goddard Space Flight Center, NASA; and P. Giommi, Satellite for Astronomy in X-rays, Scientific Data Center, communicate: “As part of a systematic search for periodicities in the ROSAT PSPC light curves (energy range 0.1–2 keV) of sources listed in the WGACAT (IAUC 6100), we discovered two new x-ray pulsators. A periodic signal at ∼13 s (significance 15σ) was detected in the x-ray flux emitted by HD 49798, a hydrogen-deficient, sdO6, single-component spectroscopic binary with an orbital period of 1.54 day. The observation was carried out on 1992 Nov. 12 with a total exposure of 5450 s. The modulation was nearly sinusoidal and energy independent, with a pulsed fraction (i.e., semi-amplitude of modulation divided by the mean source count rates) of ∼60 percent. The barycentric period was determined to be 13.1787 ± 0.0004 s. The ∼900 photons detected from the source show a very soft spectrum with a high energy excess. A blackbody-plus-power-law model provides a good fit; for column density between N_{H} = 0.4 and 1 × 10^{21} cm^{−2}, the blackbody temperature ranges between 32 and 12 eV. For the estimated distance of 650 pc, a very wide of range of unabsorbed 0.1- to 2-keV luminosity (roughly between 10^{33} and 10^{37} erg/s) is allowed. These results reveal the compact nature of the HD 49798 companion. It is unclear if this is a white dwarf or a neutron star, but a simple wind accretion model would favor the latter. Perhaps HD 49798 is the result of post high-mass x-ray binary common envelope evolution.

We detected also 8.9 s pulsations from 2E 0050.1–7247, an Einstein x-ray source in the Small Magellanic Cloud (significance of 6σ) identified with the variable B1 star AV 111 (Azzopardi and Vigneau 1982, A.Ap. Suppl. 50, 291). The source was observed at a level of ∼ 0.13 count/s (0.1–2 keV) during a 17600-s ROSAT PSPC observation during 1993 May 9–12. The signal had a barycentric period of 8.88163 ± 0.00001 s and had a nearly-sinusoidal shape with a 25-percent pulsed fraction. The source was hard (power law with photon index of 1.1 and N_{H} = 8 × 10^{20} cm^{−2}) with a luminosity (0.1–2 keV) of ∼10^{36} erg/s for a 60-kpc distance. During other ROSAT pointed observations in 1991–1992, the source was substantially fainter (factor of 20 or more) and softer. These results reveal that 2E 0050.1–7247 contains an accreting magnetic neutron star, likely in a transient Be high-mass x-ray binary. X-ray, ultraviolet, and optical monitoring of these sources is necessary.”

1995 December 21 Daniel W. E. Green